# GuardSplat: Efficient and Robust Watermarking for 3D Gaussian Splatting

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) has recently created impressive assets for various applications. However, the copyright of these assets is not well protected as existing watermarking methods are not suited for 3DGS considering security, capacity, and invisibility. Besides, these methods often require hours or even days for optimization, limiting the application scenarios. In this paper, we propose GuardSplat, an innovative and efficient framework that effectively protects the copyright of 3DGS assets. Specifically, 1) We first propose a CLIP-guided Message Decoupling Optimization module for training the message decoder, leveraging CLIP’s aligning capability and rich representations to achieve a high extraction accuracy with minimal optimization costs, presenting exceptional capability and efficiency. 2) Then, we propose a Spherical-harmonic-aware (SH-aware) Message Embedding module tailored for 3DGS, which employs a set of SH offsets to seamlessly embed the message into the SH features of each 3D Gaussian while maintaining the original 3D structure. It enables the 3DGS assets to be watermarked with minimal fidelity trade-offs and prevents malicious users from removing the messages from the model files, meeting the demands for invisibility and security. 3) We further propose an Anti-distortion Message Extraction module to improve robustness against various visual distortions. Extensive experiments demonstrate that GuardSplat outperforms the state-of-the-art methods and achieves fast optimization speed. The code will be publicly available.
@@ 1 Introduction
Figure 1: Comparisons of four 3D watermarking frameworks. These frameworks differ in how to embed messages and how to train message decoders. (a) Directly training 3D models on the watermarked images. (b) Directly embedding messages and training a message decoder for extraction. (c) Employing the message decoder from a 2D watermarking network for optimization. (d) GuardSplat first trains a message decoder to extract messages from CLIP [35] textual features. This message decoder then can be applied to the CLIP visual features for watermarking 3D models via optimization.
@@
Existing watermarking methods [60, 24, 13, 59] partially meet the four requirements above and have improved a lot for the watermark of 2D or 3D digital assets. However, they are inadequate for the 3DGS framework, which can be categorized into four groups. First, an intuitive method is directly applying 2D watermarking methods for 2D training or rendered views. For instance, employing HiDDeN [60] to watermark the 2D rendered views cannot protect the original model files. Besides, directly training 3D models on the watermarked images shown as Figure 1 (a) exhibits low bit accuracy in message extraction, because this pipeline does not guarantee novel views contain a consistent watermark. Second, some methods directly embed messages into 3D models during optimization (e.g., [24, 59], Figure 1 (b)). Although this pipeline guarantees novel views that contain a consistent watermark, it needs to optimize a new message decoder per scene during watermarking 3D models, requiring expensive optimization costs. Third, to avoid per-scene optimization, one might consider pre-training a general-purpose message decoder (e.g., [13, 12, 14], Figure 1 (c)). Given an image and a message, they first train an encoder-decoder model before finetuning the 3D representations. This encoder-decoder model aims to reconstruct the image and extract the message, where the encoder tries to keep the strict consistency between the input and output images while the decoder subsequently tries to extract the messages as intact as possible from the output image. As a result, directly using the decoder to watermark 3D models may yield degraded performance due to the fidelity-capacity trade-off. Moreover, simultaneously optimizing both encoder and decoder is time-consuming.
@@
In this paper, we present GuardSplat, a novel framework that can effectively protect the copyright of 3DGS assets. As shown in Figure 1 (d), compared to conventional 2D watermarking methods that optimize both the message encoder and decoder simultaneously, we propose a message decoupling optimization module guided by Contrastive Language-Image Pre-training (CLIP) [35], which can only train the decoder for message extraction (top row), significantly reducing the optimization costs. Thanks to the text-image aligning capability and rich representations of CLIP, this message decoder can embed large-capacity messages into 3DGS assets (bottom row) with minimal optimization costs (see Figure 2), demonstrating superior capacity and efficiency. Subsequently, we tailor a watermark embedding module for 3DGS, employing a set of spherical harmonics (SH) offsets to learn the watermarked SH features during optimization. It can seamlessly embed messages into the SH features of each 3D Gaussian while maintaining the original 3D structure, watermarking 3DGS assets with minimal trade-offs in fidelity. It also prevents malicious users from removing the watermarks from the model files, meeting the demands for invisibility and security. We further propose an anti-distortion message extraction module, which employs a differentiable distortion layer to simulate the randomly-distorted views during optimization, enabling the watermarked SH features to have strong robustness against various types of distortions. Extensive experiments on Blender [27] and LLFF [26] datasets demonstrate that our GuardSplat outperforms the state-of-the-art methods. GuardSplat achieves fast training speed, which takes 5 and 20 minutes to train the decoder and watermark a 3DGS asset on a single RTX 3090 GPU, respectively. Overall, the main contributions of this paper are summarized as follows:
  • We propose GuardSplat, a new 3D watermarking approach that can effectively protect the copyright of 3DGS assets in the real world.
@@
  • We propose a CLIP-guided Message Decoupling Optimization module to train the message decoder, achieving superior capacity and efficiency. We tailor a watermark embedding method for 3DGS to meet the invisibility and security demands. We further propose an anti-distortion message extraction for good robustness.
  • Experimental results demonstrate that our GuardSplat significantly outperforms the state-of-the-art methods and achieves fast optimization speed for training message decoder and watermarking 3DGS assets.
Figure 2: Training accuracy curves with $N_{L}=32$ bits on Blender [27] dataset. Compared to existing methods that require hours or even days for optimization, our GuardSplat achieves much better efficiency, which only takes 5 and 20 minutes to train the message decoder and watermark a 3DGS asset, respectively.
@@ 2 Related Works
3D Representations. Neural radiance field (NeRF) [27] is a compelling solution for 3D representations, which is based on the standard volumetric rendering [16] and alpha compositing techniques [33], building the implicit representations using the multi-layer perceptron (MLP). Follow-up works adapt NeRF to various domains, such as sparse-view reconstruction [54, 8, 47], acceleration [28], generative modeling [40, 29], text-to-3D generation [32, 49], anti-aliasing [3, 4], medical image super resolution [7], and RGB-D scene synthesis [1]. 3D Gaussian Splatting (3DGS) [17] has become a mainstream approach for 3D representations with its fast optimization and rendering speed. Unlike NeRF, 3DGS explicitly represents the scene using a set of 3D Gaussians and renders views through splatting [19]. It has been applied to various scenarios and created numerous assets, including text-to-3D generation [45, 52, 21], avatar generation [41, 56], single-view generation [44, 61], anti-aliasing [55], and SLAM [25, 51].
Digital Watermarking. Early studies [2, 20, 36, 46] proposed to embed the watermarks within frequency domains. With the advent of deep learning, Zhu et al. [60] proposed the first end-to-end deep watermarking framework – HiDDeN, while the subsequent advances investigate to improve the robustness [22], and extend the application scenarios [57, 23]. Recent methods [9, 50] proposed diffusion-based watermarking to protect the contents yielded from diffusion models [11, 42, 38, 39]. To protect 3D assets, most methods [30, 34] focused on embedding and detecting watermarks within meshes and point clouds. Yoo et al. [53] provided a novel perspective, which embedded the invisible watermarks into 3D models through differentiable rendering pipelines, allowing the watermarks to be extracted from rendered views. Inspired by [53], CopyRNeRF [24] and WateRF [13] aim to insert watermarks into NeRF [27]. Specifically, CopyRNeRF [24] replaces the color representations, while WateRF [13] embeds the message into model weights via optimization. Recently, Song et al. [43] propose to prevent using Triplane Gaussian Splatting (TGS) [61] for unauthorized 3D reconstruction from copyrighted images. Zhang et al. [59] introduced a steganography model for 3DGS that employs secured features to replace SH features, and trains scene and message decoders to extract views and hidden messages, respectively. Current works [12, 14] aim to directly embed messages into 3DGS models via a HiDDeN decoder. However, since HiDDeN has an inherent trade-off between fidelity and capacity, using its decoder for optimization may result in sub-optimal capacity. Moreover, [12, 14] may significantly alter the 3D structure during watermarking, leading to low-fidelity results.
Figure 3: Visual examples of CopyRNeRF [24], WateRF [13], WateRF + 3DGS, and our GuardSplat. Heatmaps at the bottom show the differences (×$\times 10$) between the watermarked and original views. Red text indicates the best performance.
@@ 3 Preliminary
Contrastive Language-Image Pre-training. Contrastive Language-Image Pre-training (CLIP) [35] is pre-trained to match images with their corresponding natural language descriptions on 400 million training image-text pairs collected from the internet. It consists of two independent encoders: a textual encoder $\mathcal{E_{T}}$ and a visual encoder $\mathcal{E_{V}}$. Given a batch of images and texts, CLIP extracts textual features $F_{\mathcal{T}}\in\mathbb{R}^{512}$ and visual features $F_{\mathcal{V}}\in\mathbb{R}^{512}$ through the textual and visual encoders, respectively. These encoders are trained to learn the aligning capability of text-image pairs by maximizing the similarity between textual and visual features via a contrastive loss.
@@ 4 Method
In this section, we propose GuardSplat to effectively protect the copyright of 3D Gaussian Splatting (3DGS) [17] assets. The overview of the proposed method is depicted in Figure 4. Specifically, we first propose a message decoupling optimization module guided by Contrastive Language-Image Pretraining (CLIP) [35] to train the message decoder $\mathcal{D_{M}}$. By analyzing 3DGS rendering pipeline, we then present a spherical-harmonic-aware (SH-aware) watermark embedding approach to integrate the watermarked SH features for a pre-trained 3DGS model. Furthermore, we design a strategy for anti-distortion message extraction. As a result, GuardSplat can not only embed large-capacity messages with minimal optimization costs, but also achieves superior fidelity and strong robustness, suggesting that GuardSplat can protect the copyright of 3DGS models without the affection of normal use.
4.1 CLIP-guided Message Decoupling Optimization
@@
To address this issue, we propose a CLIP-guided Message Decoupling Optimization module that optimizes a general-purpose message decoder and a 3DGS model, as shown in Figure 2. The key to the success of this module is that the Contrastive Language-Image Pre-training (CLIP) [35] builds a bridge between the texts and images. As discussed in Section 3, CLIP consists of a good textual encoder and a good visual encoder that is trained on a dataset of 400 million text-image pairs, providing rich text and image representations. As shown in Figure 4 (a), given a binary message $M\in\{0,1\}^{L}_{i=1}$, we first transform it into CLIP tokens $T$ using the proposed bit-to-token mapping as:
@@
where $t_{S}=49406$ and $t_{E}=49407$ denote the start and ending points of CLIP text token, respectively. $\varPhi(\cdot,\cdot)$ is a function that uniformly maps the $i$-th bit to an integer number within the range $[1,49405]$. To match the format of CLIP tokens, $T$ is then zero-padded to a size of $77$. Finally, we feed the tokens into CLIP textual encoder $\mathcal{E_{T}}$ and employ a message decoder $\mathcal{D_{M}}$ built by multi-layer perception (MLP) with 3 fully-connected (FC) layers to extract the messages $\hat{M}$ from the output textual features $F_{\mathcal{T}}\in\mathbb{R}^{512}$ as:
@@
This message decoder can be optimized by minimizing the message loss $\mathcal{L}_{\text{msg}}$ as:
@@
As a result, the message decoder can be directly optimized without being constrained by the invisibility, and thereby achieves superior capacity and training efficiency. Furthermore, CLIP’s rich representation also enables it to achieve better performance with minimal optimization costs.
4.2 Spherical-harmonic-aware Watermark Embedding
@@
Based on the above observations, we argue that it is required to maintain the original 3D structure during optimization. To achieve this, we propose an SH-aware Watermark Embedding module, a simple yet efficient approach tailored for 3DGS to watermark pre-trained models with minimal losses in fidelity. As shown in Figure 4 (b), for each 3D Gaussian, we freeze all the attributes and create a learnable SH offset $\boldsymbol{h}^{o}_{i}\in\mathbb{R}^{48}$ for watermarking. The reason behind this is the fact that SH parameters represent view-dependent effects like glossy or specular highlights, which only exist in a few regions of a scene. Thus, embedding the secret message into SH features with minimal constraints can preserve the fidelity of the 3D asset. Specifically, we first add each SH offset $\boldsymbol{h}^{o}_{i}$ to the corresponding SH coefficient $\boldsymbol{h}_{i}$ as the watermarked SH feature, and then feed it into the 3DGS rendering pipeline to render the watermarked views. To further alleviate the decline in fidelity brought by excessive offset, we employ an offset loss $\mathcal{L}_{\text{off}}$ to constrain its magnitude by:
@@
where $N$ denotes the number of 3D Gaussians. As a result, the secret message can be seamlessly embedded into the learnable SH offset of each 3D Gaussian via optimization without altering the original 3D structure, which exhibits high fidelity and prevents malicious users from removing the watermarks from the model files, achieving superior invisibility and security.
Table 1: Comparisons of the start-of-the-art methods on Blender [27] and LLFF [26] datasets for bit accuracy and reconstruction qualities w.r.t various message lengths. Bold and underline texts indicate the best and second-best performance.
@@
WateRF [13] + 3DGS 92.89 31.01 0.9678 0.0475 90.15 29.56 0.9611 0.0492 87.30 29.13 0.9562 0.0534
HiDDeN [60] + 3DGS [17] 63.07 31.59 0.9890 0.0171 52.46 34.51 0.9882 0.0209 53.08 33.42 0.9787 0.0299

### v5

@@ Abstract
3D Gaussian Splatting (3DGS) has recently created impressive 3D assets for various applications. However, considering security, capacity, invisibility, and training efficiency, the copyright of 3DGS assets is not well protected as existing watermarking methods are unsuited for its rendering pipeline. In this paper, we propose GuardSplat, an innovative and efficient framework for watermarking 3DGS assets. Specifically, 1) We propose a CLIP-guided pipeline for optimizing the message decoder with minimal costs. The key objective is to achieve high-accuracy extraction by leveraging CLIP’s aligning capability and rich representations, demonstrating exceptional capacity and efficiency. 2) We tailor a Spherical-Harmonic-aware (SH-aware) Message Embedding module for 3DGS, seamlessly embedding messages into the SH features of each 3D Gaussian while preserving the original 3D structure. This enables watermarking 3DGS assets with minimal fidelity trade-offs and prevents malicious users from removing the watermarks from the model files, meeting the demands for invisibility and security. 3) We present an Anti-distortion Message Extraction module to improve robustness against various distortions. Experiments demonstrate that GuardSplat outperforms state-of-the-art and achieves fast optimization speed.
@@ 1 Introduction
Figure 1: Comparisons of four 3D watermarking frameworks. They differ in how to embed messages and train message decoders. (a) Directly training 3D models on the watermarked images. (b) Simultaneously training a 3D model and a message decoder. (c) Employing the message decoder from a 2D watermarker for optimization. (d) GuardSplat first trains a message decoder to extract messages from CLIP [41] textual features. This message decoder then can be applied to the CLIP visual features for watermarking 3D models via optimization.
@@
Existing watermarking methods [69, 28, 14, 66] partially meet the four requirements above and have improved a lot for the watermark of 2D or 3D digital assets. However, they are inadequate for the 3DGS framework, which can be categorized into three groups. First, an intuitive method is directly applying 2D watermarking methods for 2D training or rendered views. For instance, employing HiDDeN [69] to watermark the 2D rendered views cannot protect the original model files. Besides, directly training 3D models on the watermarked images shown as Figure 1 (a) exhibits low bit accuracy in message extraction, as it cannot guarantee novel views contain a consistent watermark. Second, some methods directly embed messages into 3D models during optimization (e.g., [28, 66], Figure 1 (b)). Although it guarantees a consistent watermark across novel views, it optimizes a new message decoder per scene during watermarking 3D models, requiring expensive optimization costs. Third, to avoid per-scene optimization, one might consider a general-purpose message decoder (e.g., [14, 49, 13, 15], Figure 1 (c)), which is pre-trained from a 2D watermarking network. However, these networks are encoder-decoder models that simultaneously reconstruct the image and extract the message, where the encoder tries to keep the fidelity between the input and output images while the decoder subsequently tries to extract the messages as intact as possible from the output image. As a result, directly using the decoder to watermark 3D models may yield degraded performance due to the fidelity-capacity trade-off. Moreover, simultaneously optimizing both encoder and decoder is time-consuming.
@@
In this paper, we present GuardSplat, a novel watermarking framework to protect the copyright of 3DGS assets. As shown in Figure 1 (d), compared to conventional 2D watermarking methods that optimize both the message encoder and decoder simultaneously, we propose a message decoupling optimization module guided by Contrastive Language-Image Pre-training (CLIP) [41], which can only train the decoder for message extraction (top row), significantly reducing the optimization costs. Thanks to the text-image aligning capability and rich representations of CLIP, this decoder can embed large-capacity messages into 3DGS assets (bottom row) with minimal costs (see Figures 2 and S2), demonstrating superior capacity and efficiency. Subsequently, we tailor a message embedding module for 3DGS, employing a set of spherical harmonics (SH) offsets to seamlessly embed messages into the SH features of each 3D Gaussian while maintaining the original 3D structure. This enables watermarking 3DGS assets with minimal trade-offs in fidelity and also prevents malicious users from removing the watermarks from the model files, meeting the demands for invisibility and security. We further design an anti-distortion message extraction module, which simulates the randomly distorted views during optimization using the differentiable distortion layer, allowing the watermarked SH features to achieve strong robustness against various distortions. Extensive experiments on Blender [32] and LLFF [31] datasets demonstrate that our GuardSplat outperforms the state-of-the-art methods. GuardSplat achieves fast optimization speed, which takes 5 and 10 minutes to train the decoder and watermark a 3DGS asset on a single RTX 3090 GPU, respectively.
@@
Overall, the main contributions of this paper are summarized as follows: 1) We present GuardSplat, a new watermarking framework to protect the copyright of 3DGS assets. 2) We propose a CLIP-guided Message Decoupling Optimization module to train the message decoder, achieving superior capacity and efficiency. We tailor a message embedding method for 3DGS to meet the invisibility and security demands. We further introduce an anti-distortion message extraction for good robustness. 3) Experiments demonstrate that our GuardSplat outperforms state-of-the-art and achieves fast optimization speed for training message decoder and watermarking 3DGS assets.
Figure 2: Performance of state-of-the-art methods with $N_{L}=32$ bits on Blender [32] and LLFF [31] datasets. The radius of circles is proportional to their total training time (decoder optimization + watermarking) evaluated on RTX 3090 GPU.
@@ 2 Related Works
3D Representations. Neural radiance field (NeRF) [32] is a compelling solution for 3D representations, which is based on the standard volumetric rendering [18] and alpha compositing techniques [39], building the implicit representations using the multi-layer perceptron (MLP). Follow-up works adapt NeRF to various domains, such as sparse-view reconstruction [61, 8, 55], acceleration [33], generative modeling [46, 35], text-to-3D generation [38, 57], anti-aliasing [4, 5], medical image super-resolution [7], and RGB-D scene synthesis [2]. 3D Gaussian Splatting (3DGS) [19] has become a mainstream approach for 3D representations with its fast optimization and rendering speed. Unlike NeRF, 3DGS explicitly represents the scene using a set of 3D Gaussians and renders views through splatting [22]. It has been applied to various scenarios, including text-to-3D generation [24], avatar generation [47, 63], single-view generation [51, 71], anti-aliasing [62], and SLAM [30, 59].
Digital Watermarking. Early studies [3, 23, 42, 53, 34] proposed to embed the watermarks within frequency domains. With the advent of deep learning, Zhu et al. [69] proposed the first end-to-end deep watermarking framework – HiDDeN, while the subsequent advances investigate to improve the robustness [26, 9, 29, 1], and extend the application scenarios [64, 52, 27]. Recent methods [10, 58] proposed diffusion-based watermarking to protect the contents yielded from diffusion models [12, 48, 44, 45]. To protect 3D assets, most methods [36, 40, 67, 70] focused on embedding and detecting watermarks within meshes and point clouds. Yoo et al. [60] provided a novel perspective, which embedded the invisible watermarks into 3D models through differentiable rendering pipelines, allowing the watermarks to be extracted from rendered views. Inspired by [60], CopyRNeRF [28], WateRF [14], and NeRFProtector [49] aim to insert watermarks into NeRF [32]. Specifically, CopyRNeRF [28] replaces the color representations, while WateRF [14] and NeRFProtector [49] embed the message into model weights via optimization. Recently, Song et al. [50] propose to prevent using Triplane Gaussian Splatting (TGS) [71] for unauthorized 3D reconstruction from copyrighted images. Zhang et al. [66] introduced a steganography model for 3DGS that employs secured features to replace SH features, and trains scene and message decoders to extract views and hidden messages, respectively. Current works [13, 15] aim to directly embed messages into 3DGS models via a pre-trained 2D watermarking decoder. However, since 2D watermarking methods have an inherent trade-off between fidelity and capacity, using their decoder for optimization may result in sub-optimal capacity. Moreover, [13, 15] may significantly alter the 3D structure during watermarking, leading to low-fidelity results.
@@ 3 Preliminary
Contrastive Language-Image Pre-training. CLIP [41] is pre-trained to match images with natural language descriptions on 400 million training image-text pairs collected from the internet. It consists of two independent encoders: a textual encoder $\mathcal{E_{T}}$ and a visual encoder $\mathcal{E_{V}}$, which extract textual features $F_{\mathcal{T}}\in\mathbb{R}^{512}$ and visual features $F_{\mathcal{V}}\in\mathbb{R}^{512}$ from the given batch of images and texts, respectively. These encoders are trained to learn the aligning capability of text-image pairs by maximizing the similarity between textual and visual features via a contrastive loss.
@@ 4 Method
In this section, we propose GuardSplat to effectively protect the copyright of 3D Gaussian Splatting (3DGS) [19] assets. The overview of the proposed method is depicted in Figure 3. Specifically, we first propose a message decoupling optimization module guided by Contrastive Language-Image Pretraining (CLIP) [41] to train the message decoder $\mathcal{D_{M}}$. By analyzing the 3DGS rendering pipeline, we then present a spherical-harmonic-aware (SH-aware) message embedding module to integrate the watermarked SH features for a pre-trained 3DGS model. Furthermore, we design a strategy for anti-distortion message extraction. As a result, GuardSplat can not only embed large-capacity messages with minimal optimization costs but also achieves superior fidelity and robustness, suggesting that GuardSplat can protect the copyright of 3DGS models without the affection of normal use.
4.1 Message Decoupling Optimization
@@
To address this issue, we propose a Message Decoupling Optimization module guided by CLIP [41] that optimizes a general-purpose message decoder and a 3DGS model, as shown in Figure 1 (d). The key to the success of this module is that the Contrastive Language-Image Pre-training (CLIP) [41] builds a bridge between the texts and images. As discussed in Section 3, CLIP consists of a good textual encoder and a good visual encoder that is trained on a dataset of 400 million text-image pairs, providing rich text and image representations. As shown in Figure 3 (a), given a binary message $M\in\{0,1\}^{L}_{i=1}$, we first transform it into CLIP tokens $T$ using the proposed bit-to-token mapping as:
@@
where $t_{S}=49406$ and $t_{E}=49407$ denote the start and ending points of the CLIP text token, respectively. $\varPhi(\cdot,\cdot)$ is a function that uniformly maps the $i$-th bit to an integer number within the range $[1,49405]$. To match the format of CLIP tokens, $T$ is then zero-padded to a size of $77$. Finally, we feed the tokens into CLIP textual encoder $\mathcal{E_{T}}$ and employ a message decoder $\mathcal{D_{M}}$ built by multi-layer perception (MLP) with 3 fully-connected (FC) layers to extract the messages $\hat{M}$ from the output textual features $F_{\mathcal{T}}\in\mathbb{R}^{512}$ as:
@@
$\mathcal{D_{M}}$ can be optimized by minimizing the message loss as:
@@
As a result, the message decoder can be directly optimized without being constrained by invisibility, achieving superior capacity and training efficiency. Furthermore, CLIP’s rich representation also enables it to achieve better performance with minimal optimization costs.
4.2 SH-aware Message Embedding
@@
Based on the above observations, we argue that it is required to maintain the original 3D structure during optimization. To achieve this, we propose an SH-aware Message Embedding module, a simple yet efficient approach tailored for 3DGS to watermark pre-trained models with minimal losses in fidelity. As shown in Figure 3 (b), for each 3D Gaussian, we freeze all the attributes and create a learnable SH offset $\boldsymbol{h}^{o}_{i}\in\mathbb{R}^{48}$ for watermarking. The reason behind this is the fact that SH parameters represent view-dependent effects like glossy or specular highlights, which only exist in a few regions of a scene. Thus, embedding the secret message into SH features with minimal constraints can preserve the fidelity of the 3D asset. Specifically, we first add each SH offset $\boldsymbol{h}^{o}_{i}$ to the corresponding SH coefficient $\boldsymbol{h}_{i}$ as the watermarked SH feature, and then feed it into the 3DGS rasterization to render the watermarked views. To further alleviate the fidelity decline brought by excessive offset, we employ an offset loss to constrain its magnitude:
@@
where $N$ denotes the number of 3D Gaussians. As a result, the secret message can be seamlessly embedded into the SH offset of each 3D Gaussian in optimization, which not only maintains the original 3D structure but also prevents malicious users from removing the watermarks from the model files, achieving superior invisibility and security.
Table 1: Comparisons of the start-of-the-art methods on Blender [32] and LLFF [31] datasets for bit accuracy and reconstruction qualities w.r.t various message lengths. Bold text indicates the best performance in this table.
@@
3DGS Built on Watermarked Images
3DGS [19] + CIN [29] 56.73 31.95 0.9194 0.1027 53.13 31.74 0.9279 0.0944 55.78 30.25 0.9139 0.1120
3DGS [19] + SSL [9] 58.94 36.51 0.9737 0.0094 61.85 35.24 0.9706 0.0179 58.79 35.88 0.9710 0.0123
3DGS [19] + HiDDeN [69] 63.07 31.59 0.9790 0.0171 52.46 34.51 0.9682 0.0209 53.08 33.42 0.9687 0.0299
3DGS [19] + DwtDctSvd [34] 55.44 34.78 0.9582 0.0399 53.15 32.32 0.9477 0.0547 51.83 31.09 0.9302 0.0704
3DGS [19] + StegaStamp [52] 79.72 35.52 0.9697 0.0181 82.36 34.04 0.9601 0.0265 83.97 32.54 0.9523 0.0406
3DGS optimized by altering all attributes ($\text{Offset}_{\text{all}}$)
GaussianMarker [13] 99.36 34.42 0.9822 0.0124 98.85 33.98 0.9788 0.0163 98.25 32.12 0.9723 0.0234
3DGS [19] w/ WateRF [14] 92.89 31.01 0.9678 0.0475 90.15 29.56 0.9611 0.0492 87.30 29.13 0.9562 0.0534
GuardSplat (Ours) + 2D Watermarking Decoders
GuardSplat (Ours) + CIN [29] 95.75 37.88 0.9762 0.0092 93.35 37.42 0.9726 0.0109 92.77 37.10 0.9689 0.0124
GuardSplat (Ours) + SSL [9] 99.50 40.92 0.9935 0.0020 98.60 38.95 0.9920 0.0028 98.14 38.51 0.9909 0.0030